A certain sum of money is invested at 10%. Twice that amount is invested at 8%. The total amount of interest from both investments is $91.00. How much is invested at 8%. Can you show me how you work it out and solve it?
Answer by checkley71(8403)

.10X+.08*2X=9100
.10X+.16X=9100
.26X=9100
X=9100/.26
X=35000 AMOUNT INVESTED @ 10%
2*35000=70000 INVESTED @ 8%
PROOF
35000*.10+70000*.08=9100
3500+5600=9100
9100=9100